Bombe, Inc. produces the basic fillings used in many popular frozen desserts and treats vanilla and chocolate ice​ creams, puddings,​ meringues, and fudge.

Bombe uses standard costing and carries over no inventory from one month to the next. The​ ice-cream product​ group's results for June 2017

were as​ follows:

Performance Report, June 2017

Actual

Static

Results

Budget

Units (pounds)

360,000

348,000

Revenues

$1,839,600

$1,809,600

Variable manufacturing costs

1,400,400

1,322,400

Contribution margin

$439,200

$487,200

Tom Finn​, the business manager for​ ice-cream products, is pleased that more pounds of ice cream were sold than budgeted and that revenues were up.​ Unfortunately, variable manufacturing costs went​ up, too. The bottom line is that contribution margin declined by $ 48,000, which is less than 3​% of the budgeted revenues of $ 1,809,600.

​Overall,Finn feels that the business is running fine.

1.

Calculate the​ static-budget variance in​ units, revenues, variable manufacturing​ costs, and contribution margin. What percentage is each​ static-budget variance relative to its​ static-budget amount?

2.

Break down each​ static-budget variance into a​ flexible-budget variance and a​ sales-volume variance.

3.

Calculate the​ selling-price variance.

4.

Assume the role of management accountant at Bombe, How would you present the results to Tom Finn? Should he be more concerned? If so, why?
Answer #1

1.Static Budget Variance:

Static Budget is the Budget that does not change with the variance in the level of activity nor with the time.

Static Budget Variance is the Variance/ difference between the budgeted and actual results.

S.no Particulars Actual Results (a) Static Budget (b)

Static Budget Variance

( c = a-b )

Static Budget Variance

in %

( d = c/b*100)

1 Units ( Pounds) 360,000 348,000 12,000 3.45%
2 Revenues 1,839,600 1,809,600 30,000 1.66%
3 Variable manufacturing Cost 1,400,400 1,322,400 78,000 5.90%
4. Contribution Margin 439,200 487200 (48,000) 9.85%

We have calculated the static budget Variance and its relative percentage to the static budget amount in the above table.

First Take the static Budgeted (Column a) and actual results (Column b) and the difference between the both will give you the Static budgeted Variance which we have done in column c.

To find the relative percentage of static budget variance to the Static budget, take the static budget variance which we have calculated in the earlier step divide it with the static budget amount and multiply with 100 to get the percentage variance which we have done in column d.

2. Flexible Budget Variance and Sales Volume Variance:

Flexible Budget Variance

Flexible Budget is the budget which is changed from time to time according to the level of activity. unlike static budget which remain constant over a period of time flexible budget can be altered according to the changes in the level of activity.

here in the above question the budget is drawn for 348,000 units but the actual sales were 360,000 units. so lets prepare a flexible budget for 360,000 units and compare it with actual results to find out if there is any variance.

Particulars Actual Results (a) Static Budget (b)

Static Budget

Per unit (c)

Flexible Budget (d)

(per unit rate * actual units sold)

Flexible Budget Variance (e) = (a-d)
Units 360,000 348,000 - 360,000 -
Revenues 1,839,600 1,809,600 5.2 1,872,000 32,400
Variable Manufacturing Cost 1,400,400 1,322,400 3.8 1,368,000 32,400
Contribution Margin 439,200 487,200 1.4 504,000 64,800

First we have taken the actual results and static budget amount from the question.

Then we are going to draw flexible budget (i.e) a budget for 360,000 units and not 348,000 units as in static budget.

Lets calculate the budget cost per unit from static budget by dividing the revenues, variable manufacturing cost and contribution margin of static budget by 348,000 units ( static budget sales units). which we have done in column c

Now lets calculate Flexible budget amounts in order to do that lets take the no.of units sold as 360,000 pounds and multiply it with cost/revenue per unit which we have calculated in Column c. Now in column d we have Flexible budget amounts.

To calculate Flexible budget Variance, Subtract the actual results from Flexible budget amounts.

Sales Volume Variance

Sales volume variance is difference between actual and budgeted no.of units sold multiplied by the budgeted selling price per unit.

Sales volume variance= (Actual no. of units sold - Budgeted no. of units sold) x budgeted selling price per unit.

Here, Actual sales = 360,000 pounds

Budgeted sales = 348,000 pounds

Budgeted selling price = 1,809,600 / 348,000 = $5.2

Sales Volume Variance = (360,000 - 348,000) * 5.2 = $62,400

3.Selling Price Variance:

Selling price variance is difference between actual and budgeted selling price multiplied by the actual no. of units sold.

Selling price variance = (Actual price - Budgeted price) x Actual unit sales

Here, Actual Price  = Actual revenue / actual no. of units sold = 1,839,600 / 360,000 = $ 5.11

Budgeted Price = Budgeted revenue / Budgeted no. of units sold = 1,809,600 / 348,000 = $ 5.2

Selling price variance = (5.11 - 5.2 ) * 360,000 = $(32,400)

4. Suggestions:

As a management accountant at Bombe, I would suggest the following to Tom Finn

By calculating the above budget and variance it is found out that though the sales volume is increased to 360,000 units from budgeted 348,000 units the selling price is decreased to $5.11 from budgeted $5.2. If the sale of 360,000 units were made at $5.2 then the company would have made 1,872,000 instead of 1,839,600. so by selling at lesser price the company has lost $ 32,400 ( from flexible budget revenue row and selling price variance) . In order to increase the revenue and contribution, Selling price of the units should be increased.

By looking at the variable manufacturing cost and Flexible budget variance of the variable cost it is clear that the actual cost is more than the budgeted cost by $32,400. Measures has to be taken that the variable cost has to be reduced atleast to the budgeted level so that the contribution margin can be increased.